# Deep learning joint extremes of metocean variables using the SPAR model

### v2

@@ 3.2 Conditional radial modelling
Viewing the angular variable $\mathbf{W}$ as a ‘covariate’ for the radial variable $R$, inference for the SPAR model is analogous to a non-stationary univariate POT analysis, for which many parametric and semi-parametric approaches have been proposed [chavez2005, Randell2016, Youngman2019, Zanini2020, Barlow2023]. Non-stationary POT modelling can be performed using generalised Pareto (GP) regression (i.e., modelling GP parameters as functions of covariates), which proceeds by first estimating the threshold function, $u(\mathbf{w})$, via a quantile regression procedure, and then estimating the GP scale and shape parameter functions, $\sigma(\mathbf{w})$ and $\xi(\mathbf{w})$ respectively, via likelihood-based inference procedures. The choices of the functional forms for $u(\mathbf{w})$, $\sigma(\mathbf{w})$, and $\xi(\mathbf{w})$ determine the flexibility of the overall model. As the dimension $d$ grows, these mappings become increasingly complex, and models that represent the functions via semi-parametric models (such as the splines used in initial work with the SPAR model [murphy2024inference] and other similar angular-radial approaches [simpson2024estimating, majumder2023semiparametric]) become increasingly computationally-demanding to estimate. Consequently, we adopt a deep learning approach, whereby the threshold and GP parameter functions are represented using artificial neural networks (ANN). For details on GP regression with deep learning methods, see [Richards2024].
Although the representation of the GP parameter functions via ANNs differs from previous approaches in the SPAR framework, the ‘loss function’ used for optimisation of the model is the same. That is, we estimate the threshold and GP parameter functions that minimise the relevant loss functions evaluated on a hold-out dataset (in the case of the GP parameters, the loss function is the negative log-likelihood), as described in Section 3.2.2. This work builds upon the approach of [murphy2024deep], who use deep learning to estimate the extremal dependence structure of random vectors on standardised marginal scales via a similar angular-radial decomposition. In contrast, our approach does not require marginal transformation, and is thus not subject to marginal estimation uncertainty.
@@ 3.2.1 Neural network representation of conditional radial parameters
Several recent approaches have used deep learning for GP regression; see, e.g., [wilson2022deepgpd, richards2023insights, cisneros2024deep, pasche2024neural]. In these approaches, ANNs are used to model the relationships between covariates and GP parameter and threshold functions. Our approach is analogous, with the covariates taken to be angles on the hypersphere. Note that the hypersphere is compact (i.e., closed and bounded), and that this is a desirable property for extrapolation in deep learning.
Here, we design two neural network models; one for the threshold function $u(\mathbf{w})$ and one for the parameter vector $(\nu(\mathbf{w}),\xi(\mathbf{w}))$, where $\nu(\mathbf{w})=\sigma(\mathbf{w})(\xi(\mathbf{w})+1)$ is the modified scale parameter. Unlike $\sigma(\mathbf{w})$, the modified scale parameter $\nu(\mathbf{w})$ is orthogonal to $\xi(\mathbf{w})$, which helps to mitigate numerical instabilities during model fitting [cox1987parameter, chavez2005, tendijck2024practical]; [pasche2024neural] show that this is particularly helpful when fitting deep GP regression models. Both $u(\mathbf{w})$ and $(\nu(\mathbf{w}),\xi(\mathbf{w}))$ are modelled by multi-layer perceptrons (MLPs), which are a standard class of fully-connected ANN that compose multiple layers of ‘neurons’ [goodfellow2016deep]. Each neuron passes a linear combination of input variables through a nonlinear ‘activation function’, and the output is then passed to the subsequent layer; detailed discussions and illustrative figures can be found in [Richards2024], and an example schematic for an MLP representation of $(\nu(\mathbf{w}),\xi(\mathbf{w}))$ is presented in Figure 2. Inference for the parameter functions involves estimating the linear coefficients (the ‘weights’ and ‘biases’) in each neuron of the corresponding MLP.
@@ 3.2.2 Estimating the neural network parameters
To obtain estimates of the MLP parameter sets, $\mathcal{W}_{u}$ and $\mathcal{W}_{(\nu,\xi)}$, we optimise specified loss functions. Suppose that we have a set of radial and angular observations $\{(r_{i},\mathbf{w}_{i});i=1,...,n\}$. Recall from Section 2.1 that the threshold $u(\mathbf{w})$ is taken to be the quantile of $R|(\mathbf{W}=\mathbf{w})$ at exceedance probability $\zeta$. We thus can estimate $u(\mathbf{w})$ (and its corresponding parameter set $\mathcal{W}_{u}$) using quantile regression techniques [koenker2017handbook]. In this case, an appropriate loss function for $u(\mathbf{w})$ is the tilted loss,
@@
After estimation of $u(\mathbf{w})$, we define $I_{u}:=\{i\in\{1,...,n\}:r_{i}>u(\mathbf{w}_{i})\}$ as the set of indices of radial threshold exceedances. The MLP that defines the GP parameter functions can be considered as a ‘conditional density estimation network’, with the negative log-likelihood function used for optimisation; see, e.g., [neuneier1994estimation, rothfuss2019conditional]. In this case, we perform maximum likelihood estimation, with loss
@@
Optimisation of both losses, (13) and (14), proceeds via stochastic gradient descent and the ADAM algorithm [kingma2014adam]. To mitigate overfitting, data are split into training (80%) and validation (20%) sets, with the latter used to check for parameter convergence. We refer the reader to [murphy2024deep] for a more detailed overview of the fitting procedure.
@@
As noted in Section 2, when the GP shape parameter $\xi(\mathbf{w})$ is negative, the distribution of $R|(\mathbf{W}=\mathbf{w})$ has a finite upper endpoint. Training of a deep GP regression model which permits negative shape parameter values can be computationally troublesome; see discussion by [richards2022regression]. At a given angle $\mathbf{w}_{i}$, if $\xi(\mathbf{w}_{i})<0$ and the radial observation $r_{i}$ exceeds the upper endpoint, i.e., $r_{i}>u(\mathbf{w}_{i})-\sigma(\mathbf{w}_{i})/\xi(\mathbf{w}_{i})$, then the loss function in (14) will evaluate to a non-finite value. Consequently, the loss surface over which we optimise $\mathcal{W}_{(\nu,\xi)}$ is highly irregular, and iterative gradient descent methods (like ADAM) may have trouble finding global maxima, or may predict out-of-sample parameter estimates that are infeasible, i.e., the loss is non-finite. To circumvent these issues during training, we initialise the MLP to ensure that the shape parameter function $\xi(\mathbf{w})$ is non-negative for all angles $\mathbf{w}$; in this way, at the outset of the training procedure, the loss function is guaranteed to be finite for all $w_{i},i=1,\dots,n$. Then, if the gradient descent optimisation algorithm produces non-finite loss values during training, we restart training (from the last iteration with finite loss values) with a smaller learning rate. In our application we found that the fully-trained MLP provided negative values of $\xi(\mathbf{w})$ for all $\mathbf{w}\in\mathbb{S}^{d-1}$. We note that the issue of non-finite loss occurs in other likelihood-based inference procedures for GP regression, and is not a problem specific to MLPs.
3.2.3 Selecting an architecture
@@
An important choice when fitting a neural network model is the choice of architecture: this corresponds to the set of hyperparameters introduced in Section 3.2.1. There is no ‘best practice’ for this selection within the deep GP regression literature [Richards2024], and the appropriate architecture is likely to be domain specific; see [goodfellow2016deep]. Selecting a model with more hidden layers and more neurons results in higher flexibility, but at the cost of increased parameter variability and computational expense. In the spirit of parsimony, we wish to select the simplest model possible while still capturing the observed variability in the threshold and GP parameter functions over the angular domain.
@@
To select our ‘optimal’ architecture for the application detailed in Section 4, we perform a grid-search over architecture choices. For each configuration, we estimate a range of model fit diagnostics; these are discussed below. The optimal architecture is then chosen as the configuration which visually provides the best model diagnostics. We found that, for both MLP models, a simple architecture is preferable: $L=3$ hidden layers, with $h_{l}=16$ neurons per layer. This results in two MLPs, each comprising approximately 650 estimable parameters; inference for these models is not computationally demanding, and can be conducted on a standard laptop.
@@
As with univariate POT models, selecting a suitable threshold for our model is critical. Too low a threshold will result in the asymptotic arguments motivating the use of the GP model not being applicable, causing bias; whereas too high a threshold will result in too few observations for fitting, resulting in higher variance. The process of threshold selection for our model is directly analogous to that in univariate problems. That is, we fit the model for a range of threshold exceedance probabilities $\zeta$, and check for stability of inferences and goodness of fit. The threshold is then selected as the largest value of $\zeta$ for which inferences are approximately stable for $\zeta_{0}<\zeta$. In our application, we found that $\zeta=0.1$ was suitable.
@@ 4.1 Dataset
In this section, we consider the application of the SPAR model to a hindcast dataset consisting of 31-years of wind and wave variables from 01/01/1990 to 31/12/2020, for a site in the Celtic Sea, off the south-west coast of the UK, which has been identified for development of floating wind farm projects. The dataset consists of hourly values of significant wave height $(H_{s})$, mean wave period $(T_{m})$, mean wave direction $(\theta_{wave})$, hourly mean wind speed at 10 m above sea level $(U_{10})$, and wind direction $(\theta_{wind})$. Both wind and wave directions are defined as the directions in which the winds or waves are ‘going to’. These variables all influence the motion and loading of floating wind turbines, and understanding their joint extremes is important for design. Since directional variables are periodic, it does not make sense to talk about ‘extreme directions’. Instead, we work with the x- and y-components of wave height and wind speed, defined as $H_{x}=H_{s}\cos(\theta_{wave})$, $H_{y}=H_{s}\sin(\theta_{wave})$, $U_{x}=U_{10}\cos(\theta_{wind})$, and $U_{y}=U_{10}\sin(\theta_{wind})$.
@@
Unlike many classical modelling approaches, the SPAR approach extrapolates in all directions, allowing one to perform inference in any extreme region of interest. When variables have a defined lower bound at zero, as is the case for many physical quantities, the SPAR model should be able to infer this directly from the data, and these physical limits should correspond to the upper bounds of the GP model for the radial variable at the relevant angles. However, inferences at endpoints are highly uncertain since they correspond to zero exceedance probability, and consequently the model may infer a slightly negative bound in some directions. It is therefore safer to work with variables that do not have hard lower bounds at zero, so that uncertainties in the radial endpoint do not result in estimates that are not physically possible. The directional variables $H_{x}$, $H_{y}$, $U_{x}$, and $U_{y}$ are all defined on (,)($\infty$,$\infty$)( - ∞ , ∞ ), although clearly they will have some upper and lower bounds due to physical constraints (in general, we would expect the region of variable space in which the density is non-zero to be bounded, due to physical limits). For the period variable, we address the problem of the lower bound by defining $L_{T}=\log(T_{m})\in(-\infty,\infty)$ and using this variable in the model.
@@ 4.2 Normalisation and choice of origin
As different physical variables have different scales, we normalise each variable by its standard deviation. Without this, angles would tend to be clustered in the plane of whichever variable has the largest scale (wind speed in this case). We also need to define an origin in order to transform to polar coordinates. In previous work using SPAR [mackay2025spar, murphy2024inference], the origin was defined at the mean of each variable. In our application, more care must be taken when defining an appropriate origin. For the model to provide a useful description of the extremes at all angles, the support of the density, $\mathrm{supp}(f_{\mathbf{X}}):=\{\mathbf{x}\in\mathbb{R}^{d}:f_{\mathbf{X}}(\mathbf{x})>0\}$, must be star-shaped with respect to the chosen origin [hansen2020]. That is, given an origin $\mathbf{x}_{0}\in\mathrm{supp}(f_{\mathbf{X}})$ and any point $\mathbf{x}\in\mathrm{supp}(f_{\mathbf{X}})$, the line segment from $\mathbf{x}_{0}$ to $\mathbf{x}$ is contained in $\mathrm{supp}(f_{\mathbf{X}})$. Under this assumption, all rays from the origin reach the ‘edges’ of the distribution without passing through regions which have zero density. In this way, the data-cloud has a well-defined ‘inside’ and ‘outside’, with the ‘outer’ region considered ‘extreme’, and the GP-based representation of the radial component in this region is reasonable. This assumption can be verified by checking plots of the density of observations along various rays from the origin, e.g., histograms of the observed radial variable within small angular ranges. However, as discussed further below, in higher dimensional spaces, a large number of angles are required to obtain a reasonable coverage of the surface of $\mathbb{S}^{d-1}$, and visual inspection of plots of the radial density over each angular range is time consuming. Ultimately a data-driven approach for selecting an optimal choice of origin would be best.
@@
In this example, we use knowledge of the physical processes to define an appropriate origin. Firstly, wave breaking limits the maximum possible wave height for a given wave period, with the limit related to wave steepness given by $s=2\pi H_{s}/(gT_{m}^{2})$. In the three-dimensional subspace containing the variables $(H_{x},H_{y},T_{m})$, this results in a conical-shaped bound on the data, centred along the axis $H_{x}=H_{y}=0$, with the radius of the cone given by $H_{s,max}=s_{max}gT_{m}^{2}/(2\pi)$, where $s_{max}$ is the limiting steepness. The value of $s_{max}$ depends on the water depth and wind speed (among other factors), but the maximum value was found to be around $s_{max}\approx 0.08$ for our dataset. This conical shape to the distribution (see Figures 3 and 7) suggests that an appropriate choice of origin should be somewhere on the axis $H_{x}=H_{y}=0$. Due to the physical dependence of wave height on wind speed, we also locate the origin at $U_{x}=U_{y}=0$. The choice of origin for $L_{T}$ is somewhat arbitrary, but experimentation showed that using the mean value of $L_{T}$ gave satisfactory results.

### v3

@@ 3.2 Conditional radial modelling
Viewing the angular variable $\mathbf{W}$ as a ‘covariate’ for the radial variable $R$, inference for the SPAR model is analogous to a non-stationary univariate POT analysis, for which many parametric and semi-parametric approaches have been proposed [42, 43, 44, 45, 46]. Non-stationary POT modelling can be performed using generalised Pareto (GP) regression (i.e., modelling GP parameters as functions of covariates), which proceeds by first estimating the threshold function, $u(\mathbf{w})$, via a quantile regression procedure, and then estimating the GP scale and shape parameter functions, $\sigma(\mathbf{w})$ and $\xi(\mathbf{w})$ respectively, via likelihood-based inference procedures. The choices of the functional forms for $u(\mathbf{w})$, $\sigma(\mathbf{w})$, and $\xi(\mathbf{w})$ determine the flexibility of the overall model. As the dimension $d$ grows, these mappings become increasingly complex, and models that represent the functions via semi-parametric models (such as the splines used in initial work with the SPAR model [23] and other similar angular-radial approaches [47, 48]) become increasingly computationally-demanding to estimate. Consequently, we adopt a deep learning approach, whereby the threshold and GP parameter functions are represented using artificial neural networks (ANN). For details on GP regression with deep learning methods, see [49].
Although the representation of the GP parameter functions via ANNs differs from previous approaches in the SPAR framework, the ‘loss function’ used for optimisation of the model is the same. That is, we estimate the threshold and GP parameter functions that minimise the relevant loss functions evaluated on a hold-out dataset (in the case of the GP parameters, the loss function is the negative log-likelihood), as described in Section 3.2.2. This work builds upon the approach of [50], who use deep learning to estimate the extremal dependence structure of random vectors on standardised marginal scales via a similar angular-radial decomposition. In contrast, our approach does not require marginal transformation, and is thus not subject to marginal estimation uncertainty.
@@ 3.2.1 Neural network representation of conditional radial parameters
Several recent approaches have used deep learning for GP regression; see, e.g., [51, 52, 53, 54]. In these approaches, ANNs are used to model the relationships between covariates and GP parameter and threshold functions. Our approach is analogous, with the covariates taken to be angles on the hypersphere. Note that the hypersphere is compact (i.e., closed and bounded), and that this is a desirable property for extrapolation in deep learning.
Here, we design two neural network models; one for the threshold function $u(\mathbf{w})$ and one for the parameter vector $(\nu(\mathbf{w}),\xi(\mathbf{w}))$, where $\nu(\mathbf{w})=\sigma(\mathbf{w})(\xi(\mathbf{w})+1)$ is the modified scale parameter. Unlike $\sigma(\mathbf{w})$, the modified scale parameter $\nu(\mathbf{w})$ is orthogonal to $\xi(\mathbf{w})$, which helps to mitigate numerical instabilities during model fitting [55, 42, 56]; [54] show that this is particularly helpful when fitting deep GP regression models. Both $u(\mathbf{w})$ and $(\nu(\mathbf{w}),\xi(\mathbf{w}))$ are modelled by multi-layer perceptrons (MLPs), which are a standard class of fully-connected ANN that compose multiple layers of ‘neurons’ [57]. Each neuron passes a linear combination of input variables through a nonlinear ‘activation function’, and the output is then passed to the subsequent layer; detailed discussions and illustrative figures can be found in [49], and an example schematic for an MLP representation of $(\nu(\mathbf{w}),\xi(\mathbf{w}))$ is presented in Figure 2. Inference for the parameter functions involves estimating the linear coefficients (the ‘weights’ and ‘biases’) in each neuron of the corresponding MLP.
@@ 3.2.2 Estimating the neural network parameters
To obtain estimates of the MLP parameter sets, $\mathcal{W}_{u}$ and $\mathcal{W}_{(\nu,\xi)}$, we optimise specified loss functions. Suppose that we have a set of radial and angular observations $\{(r_{i},\mathbf{w}_{i});i=1,...,n\}$. Recall from Section 2.1 that the threshold $u(\mathbf{w})$ is taken to be the quantile of $R|(\mathbf{W}=\mathbf{w})$ at exceedance probability $\zeta$. We thus can estimate $u(\mathbf{w})$ (and its corresponding parameter set $\mathcal{W}_{u}$) using quantile regression techniques [58]. In this case, an appropriate loss function for $u(\mathbf{w})$ is the tilted loss,
@@
After estimation of $u(\mathbf{w})$, we define $I_{u}:=\{i\in\{1,...,n\}:r_{i}>u(\mathbf{w}_{i})\}$ as the set of indices of radial threshold exceedances. The MLP that defines the GP parameter functions can be considered as a ‘conditional density estimation network’, with the negative log-likelihood function used for optimisation; see, e.g., [59, 60]. In this case, we perform maximum likelihood estimation, with loss
@@
We first train the ANN for the threshold, and use this to define the set of threshold exceedances. The ANN for the GP parameters is trained afterwards. Optimisation of both losses, (13) and (14), proceeds via stochastic gradient descent and the ADAM algorithm [61]. To mitigate overfitting, data are split into training (80%) and validation (20%) sets, with the latter used to check for parameter convergence. We refer the reader to [50] for a more detailed overview of the fitting procedure.
@@
As noted in Section 2, when the GP shape parameter $\xi(\mathbf{w})$ is negative, the distribution of $R|(\mathbf{W}=\mathbf{w})$ has a finite upper endpoint. Training of a deep GP regression model which permits negative shape parameter values can be computationally troublesome; see discussion by [62]. At a given angle $\mathbf{w}_{i}$, if $\xi(\mathbf{w}_{i})<0$ and the radial observation $r_{i}$ exceeds the upper endpoint, i.e., $r_{i}>u(\mathbf{w}_{i})-\sigma(\mathbf{w}_{i})/\xi(\mathbf{w}_{i})$, then the loss function in (14) will evaluate to a infinite value. Consequently, the loss surface over which we optimise $\mathcal{W}_{(\nu,\xi)}$ is highly irregular, and iterative gradient descent methods (like ADAM) may have trouble finding global maxima, or may predict out-of-sample parameter estimates that are infeasible, i.e., the loss is infinite. To circumvent these issues during training, we initialise the MLP to ensure that the shape parameter function $\xi(\mathbf{w})$ is non-negative for all angles $\mathbf{w}$; in this way, at the outset of the training procedure, the loss function is guaranteed to be finite for all $w_{i},i=1,\dots,n$. Then, if the gradient descent optimisation algorithm produces infinite loss values during training, we restart training (from the last iteration with finite loss values) with a smaller learning rate. In our application we found that the fully-trained MLP provided negative values of $\xi(\mathbf{w})$ for all $\mathbf{w}\in\mathbb{S}^{d-1}$. We note that the issue of infinite loss occurs in other likelihood-based inference procedures for GP regression, and is not a problem specific to MLPs.
3.2.3 Selection of hyperparameters
@@
The two key choices for the proposed inference scheme are the architecture of the neural network and the threshold non-exceedance probability, $\zeta$. There is no ‘best practice’ for this selection within the deep GP regression literature [49], and the appropriate architecture is likely to be domain specific; see [57]. The neural network architecture determines the flexibility of the model. Selecting a model with more hidden layers and more neurons results in higher flexibility, but at the cost of increased parameter variability and computational expense. Moreover, the optimal architecture is intrinsically linked to the optimal $\zeta$, as having more exceedances, i.e., more data, allows for estimation of a more flexible model. In the spirit of parsimony, we wish to select the simplest model possible while still capturing the observed variability in the threshold and GP parameter functions over the angular domain.
@@
To select our ‘optimal’ architecture and $\zeta$ for the application detailed in Section 4, we perform a grid-search over architecture choices and a sequence of threshold exceedance probabilities in the range $0.0125\leq\zeta\leq 0.25$ at intervals of $0.0125$. For each configuration, we examine a range of goodness-of-fit diagnostic plots, discussed in Section 4.5. The optimal architecture and $\zeta$ is then chosen as those values which visually provide the best model diagnostics. For both the threshold $u(\mathbf{w})$ and GP excess models, a simple neural network architecture is preferable: $L=3$ hidden layers, with $h_{l}=16$ neurons per layer. This results in two MLPs, each comprising approximately 650 estimable parameters; inference for these models is not computationally demanding, and can be conducted on a standard laptop
@@
As with univariate POT models, selecting a suitable threshold for our model is critical. Too low a threshold will result in the asymptotic arguments motivating the use of the GP model not being applicable, causing bias; whereas too high a threshold will result in too few observations for fitting, resulting in higher variance. Thus, alongside the aforementioned visual diagnostics, we further justify the choice of $\zeta$ using a process of threshold selection which is directly analogous to that in univariate problems. For a fixed choice of architecture (as given above), we fit the model for a sequence of $\zeta$ values and check for stability of estimates of the GP shape parameter and conditional radial quantiles. The optimal exceedance probability is then selected as the smallest value of $\zeta$ for which inferences are approximately stable for $\zeta_{0}>\zeta$ (allowing for sampling uncertainty). As the GP shape parameter and radial quantiles are non-stationary, we need to assess how their values vary with threshold for a range of different angles. To give a diagnostic that is weighted towards regions with higher angular density, we consider the distribution of shape parameter and quantile estimates over all observed angles. The information is summarised as box plots, as shown in Figure 3. In our application, we found that $\zeta=0.1$ gave the best results.
@@
Figure 3: Box-plots of GP shape parameter estimates (left) and conditional radial quantile at exceedance level $10^{-6}$ (right) at observed angles for various threshold non-exceedance probabilities, $1-\zeta$.
@@ 4.1 Dataset
In this section, we consider the application of the SPAR model to a hindcast dataset consisting of 31-years of wind and wave variables from 01/01/1990 to 31/12/2020, for a site in the Celtic Sea, off the south-west coast of the UK, which has been identified for development of floating wind farm projects. The data was produced using a one-way coupled wave-current model, described in [63]. At the site considered, the mean water depth is 53 m with a 7 m tidal range. The dataset consists of hourly values of significant wave height $(H_{s})$, mean wave period $(T_{m})$, mean wave direction $(\theta_{wave})$, hourly mean wind speed at 10 m above sea level $(U_{10})$, and wind direction $(\theta_{wind})$. Both wind and wave directions are defined as the directions in which the winds or waves are ‘going to’, and directions are defined clockwise from North. These variables all influence the motion and loading of floating wind turbines, and understanding their joint extremes is important for design. Since directional variables are periodic, it does not make sense to talk about ‘extreme directions’. Instead, we work with the x- and y-components of wave height and wind speed, defined as $H_{x}=H_{s}\cos(\theta_{wave})$, $H_{y}=H_{s}\sin(\theta_{wave})$, $U_{x}=U_{10}\cos(\theta_{wind})$, and $U_{y}=U_{10}\sin(\theta_{wind})$.
@@
Unlike many classical modelling approaches, the SPAR approach extrapolates in all directions, allowing one to perform inference in any extreme region of interest. When variables have a defined lower bound at zero, as is the case for many physical quantities, the SPAR model should be able to infer this directly from the data, and these physical limits should correspond to the upper bounds of the GP model for the radial variable at the relevant angles. However, inferences at endpoints are highly uncertain since they correspond to zero exceedance probability, and consequently the model may infer a slightly negative bound in some directions. It is therefore safer to work with variables that do not have hard lower bounds at zero, so that uncertainties in the radial endpoint do not result in estimates that are not physically possible. The directional variables $H_{x}$, $H_{y}$, $U_{x}$, and $U_{y}$ are all defined on $(-\infty,\infty)$, although clearly they will have some upper and lower bounds due to physical constraints (in general, we would expect the region of variable space in which the density is non-zero to be bounded, due to physical limits). For the period variable, we address the problem of the lower bound by defining $L_{T}=\log(T_{m})\in(-\infty,\infty)$ and using this variable in the model.
@@ 4.2 Normalisation and choice of origin
As different physical variables have different scales, we normalise each variable by its standard deviation. Without this, angles would tend to be clustered in the plane of whichever variable has the largest scale (wind speed in this case). We also need to define an origin in order to transform to polar coordinates. In previous work using SPAR [24, 23], the origin was defined at the mean of each variable. In our application, more care must be taken when defining an appropriate origin. For the model to provide a useful description of the extremes at all angles, the support of the density, $\mathrm{supp}(f_{\mathbf{X}}):=\{\mathbf{x}\in\mathbb{R}^{d}:f_{\mathbf{X}}(\mathbf{x})>0\}$, must be star-shaped with respect to the chosen origin [64]. That is, given an origin $\mathbf{x}_{0}\in\mathrm{supp}(f_{\mathbf{X}})$ and any point $\mathbf{x}\in\mathrm{supp}(f_{\mathbf{X}})$, the line segment from $\mathbf{x}_{0}$ to $\mathbf{x}$ is contained in $\mathrm{supp}(f_{\mathbf{X}})$. Under this assumption, all rays from the origin reach the ‘edges’ of the distribution without passing through regions which have zero density. In this way, the data-cloud has a well-defined ‘inside’ and ‘outside’, with the ‘outer’ region considered ‘extreme’, and the GP-based representation of the radial component in this region is reasonable. This assumption can be verified by checking plots of the density of observations along various rays from the origin, e.g., histograms of the observed radial variable within small angular ranges. However, as discussed further below, in higher dimensional spaces, a large number of angles are required to obtain a reasonable coverage of the surface of $\mathbb{S}^{d-1}$, and visual inspection of plots of the radial density over each angular range is time consuming. Ultimately a data-driven approach for selecting an optimal choice of origin would be preferable.
@@
In this example, we use knowledge of the physical processes to define an appropriate origin. Firstly, wave breaking limits the maximum possible wave height for a given wave period, with the limit related to wave steepness given by $s=2\pi H_{s}/(gT_{m}^{2})$. In the three-dimensional subspace containing the variables $(H_{x},H_{y},T_{m})$, this results in a conical-shaped bound on the data, centred along the axis $H_{x}=H_{y}=0$, with the radius of the cone given by $H_{s,max}=s_{max}gT_{m}^{2}/(2\pi)$, where $s_{max}$ is the limiting steepness. The value of $s_{max}$ depends on the water depth and wind speed (among other factors), but the maximum value was found to be around $s_{max}\approx 0.08$ for our dataset (consistent with values in offshore standards, e.g. [65]) . This conical shape to the distribution (see Figures 4 and 8) suggests that an appropriate choice of origin should be somewhere on the axis $H_{x}=H_{y}=0$. Due to the physical dependence of wave height on wind speed, we also locate the origin at $U_{x}=U_{y}=0$. The choice of origin for $L_{T}$ is somewhat arbitrary, but experimentation showed that using the mean value of $L_{T}$ gave satisfactory results.